OpenStudy (anonymous):

If AM between pth and qth terms of an AP be equal to the AM between rth and sth terms of the AP then prove that p+q =r +s

Parth (parthkohli):

lol, simple proof

Parth (parthkohli):

what is the expression for AM between two numbers?

OpenStudy (anonymous):

Sorry I managed prove lol , it was simple

OpenStudy (anonymous):

It is not that formula , chuck that

Parth (parthkohli):

...?\[\dfrac{a_p + a_q}{2}=\dfrac{a_{r} + a_s }{2}\]\[\implies a_p + a_q = a_r + a_s\]\[\implies a + (p-1)d = a + (q-1)d = a + (r-1)d + a + (s-1)d\]Yup.

OpenStudy (anonymous):

Yes yes , all will cancel out and p+q=r+s Since i was doing n means , i read it as AM's

Parth (parthkohli):

OK, no problem! I studied sequences and series a few weeks ago at coaching, so feel free to tag me. I haven't practiced this for a few days. :)

OpenStudy (anonymous):

Ok , sure